# Central and Central-Parietal EEG Signatures of Parkinson’s Disease

### v1

@@ Abstract
This study investigates EEG as a potential early biomarker by applying deep learning techniques to resting-state EEG recordings from 31 subjects (15 with PD and 16 healthy controls). EEG signals were rigorously preprocessed to remove tremor artifacts, then converted to wavelet-based images by grouping spatially adjacent electrodes into triplets for convolutional neural network (CNN) classification. Our analysis across different brain regions and frequency bands showed distinct spatial-spectral patterns of PD-related neural oscillations. We identified high classification accuracy (74%) in the gamma band (40-62.4 Hz) for central-parietal electrodes (CP1, Pz, CP2), and 76% accuracy using central electrodes (C3, Cz, C4) with full-spectrum 0.4-62.4 Hz. In particular, we observed pronounced right-hemisphere involvement, specifcialy in parieto-occipital regions. Unlike previous studies that achieved higher accuracies by potentially including tremor artifacts, our approach isolates genuine neurophysiological alterations in cortical activity. These findings suggest that specific EEG-based oscillatory patterns, especially central-parietal gamma activity, may provide diagnostic information for PD, potentially before the onset of motor symptoms.
@@
Parkinson’s disease , Electroencephalography , Deep learning , Gamma oscillations , Parieto-occipital cortex , Biomarkers
@@ 1 Introduction
Parkinson’s disease (PD) is a progressive neurodegenerative disorder characterized primarily by motor symptoms including resting tremor, rigidity, bradykinesia, and postural instability [1]. PD is assocaited with the loss of dopaminergic neurons in the substantia nigra, resulting in dopamine deficiency [2]. Despite decades of research, current diagnostic approaches still rely heavily on clinical observation of motor symptoms and response to dopaminergic medication—a significant limitation given that these manifestations typically appear only after 60-80% of substantia nigra neurons have already degenerated [3].
@@
The spatial localization of PD-related EEG abnormalities presents another substantial challenge. Traditional emphasis on basal ganglia dysfunction in PD has led many researchers to focus predominantly on central and frontal electrodes, often neglecting posterior regions. Yet emerging evidence suggests significant involvement of parietal and occipital cortices in PD pathophysiology, with Tessitore et al. demonstrating altered resting-state functional connectivity in posterior cortical networks that correlates with specific cognitive symptoms. The reliance on predetermined regions of interest in most previous studies has potentially obscured important spatial patterns of EEG abnormalities across the cortex.
@@
The present study addresses these gaps by employing a deep learning framework that systematically evaluates EEG activity across multiple brain regions and frequency bands without preconceived notions about which combinations should prove most discriminative. By transforming EEG signals into wavelet-based images and grouping spatially adjacent electrodes, the approach captures local patterns of neural synchronization while avoiding the limitations of traditional frequency analysis methods. This methodology enables comprehensive mapping of PD-related EEG alterations across the entire cortex and frequency spectrum.
@@
    To systematically evaluate the discriminative power of different brain regions (electrode groups) and frequency bands without the regional biases evident in previous research, employing rigorous statistical controls appropriate for multiple comparisons across the electrode/frequency space.
@@
    To identify specific spatial-spectral patterns yielding optimal classification accuracy, with particular attention to previously under-investigated regions and frequency combinations that might serve as novel neurophysiological biomarkers for PD.
@@ 2.1 Dataset
It is important to note that the dataset has certain limitations acknowledged by the original data collectors. The Unified Parkinson’s Disease Rating Scale (UPDRS) assessments were conducted by laboratory personnel who had completed online training rather than by board-certified neurologists, which may introduce some uncertainty in clinical evaluations. Additionally, given the heterogeneous presentation of PD across patients and the relatively small sample size, we employed a cross-validation approach in our analysis to maximize the statistical validity of our findings while acknowledging the limitations inherent to the dataset size.
Figure 7 directly compares the gamma band topographical maps across all three window lengths (5, 10, and 20 seconds), highlighting the consistency of the central-parietal focus of high discriminative power regardless of the temporal window used for analysis.
@@ 2.2 EEG Preprocessing
The EEG data was downsampled from 512 Hz to 128 Hz using the EEGLAB function pop_resample(EEG, 128). This step reduced computational load while maintaining sufficient temporal resolution for frequency analysis up to 64 Hz.
Then, a finite impulse response (FIR) bandpass filter was applied using the function pop_firws(EEG, ’fcutoff’, [0.4 63.5],…) to remove frequency components below 0.4 Hz and above 63.5 Hz. This eliminated slow drifts, DC offset, and high-frequency noise while preserving the frequency bands of interest (delta, theta, alpha, beta, and gamma).
@@ 2.3 Data Preparation
    It increased the effective sample size by creating multiple examples from each participant’s recording.
@@
    It allowed for the comparison of classification performance across different temporal scales.
  3. 3.
    It provided a means to assess the stability and consistency of the identified EEG patterns.
For each window, we applied continuous wavelet transform (CWT) to convert the one-dimensional time series into two-dimensional time-frequency representations (scalograms). The resulting wavelet coefficients were analyzed across six frequency bands: delta (0.4-3.99 Hz), theta (4.0-7.79 Hz), alpha (7.8-15.59 Hz), beta1 (15.6-31.19 Hz), beta2 (31.2-39.99 Hz), and gamma (40-62.4 Hz).
To prepare the data for input to the convolutional neural network (CNN), we organized the EEG channels into 20 groups of three spatially adjacent electrodes based on their proximity on the scalp. Figure 1 shows 6 triplets (groups of 3 channels) showing the highest classification accuracy. Altogether these were the electrode triplets used in this study:
{AF3, Fz, AF4}, {FC1, Fz, FC2}, {C3, Cz, C4}, {FC5, F3, FC1}, {FC6, F4, FC2}
{CP1, Pz, CP2}, {P3, Pz, P4}, {PO3, Pz, PO4}, {CP5, C3, CP1}, {CP6, C4, CP2}
{O1, Oz, O2}, {PO3, O1, Oz}, {PO4, O2, Oz}, {P7, CP5, P3}, {P8, CP6, P4}
{F7, T7, P7}, {F3, C3, P3}, {Cz, Pz, Oz}, {F4, C4, P4}, {F8, T8, P8}
@@ 3.1 Overall Classification Performance
5-second windows 61.2% $\pm$ 27.8%
10-second windows 63.7% $\pm$ 28.9%
20-second windows 65.3% $\pm$ 29.6%
Table 3: Mean classification accuracy across all electrode triplets for different window lengths using the full frequency spectrum. Longer windows generally yield higher classification accuracy.
3.2 Electrode Groupings
To facilitate analysis based on functional neuroanatomy, we organized the 20 electrode triplets into four major groups based on their spatial distribution across the scalp (Table 4).
Group Electrode Triplets Brain Regions
1 {AF3, Fz, AF4}, {FC1, Fz, FC2}, {C3, Cz, C4}, Frontal and
{FC5, F3, FC1}, {FC6, F4, FC2} Fronto-Central Regions
2 {CP1, Pz, CP2}, {P3, Pz, P4}, {PO3, Pz, PO4}, Central-Parietal
{CP5, C3, CP1}, {CP6, C4, CP2} Regions
3 {O1, Oz, O2}, {PO3, O1, Oz}, {PO4, O2, Oz}, Parieto-Occipital
{P7, CP5, P3}, {P8, CP6, P4} Regions
4 {F7, T7, P7}, {F3, C3, P3}, {Cz, Pz, Oz}, Longitudinal/Cross-Regional
{F4, C4, P4}, {F8, T8, P8} Connections
@@ 3.4 Group-Level Performance
1: Frontal/Fronto-Central 64% $\pm$ 29% 66% $\pm$ 30% 69% $\pm$ 31%
2: Central-Parietal 63% $\pm$ 29% 65% $\pm$ 30% 67% $\pm$ 30%
3: Parieto-Occipital 61% $\pm$ 28% 64% $\pm$ 29% 68% $\pm$ 31%
4: Longitudinal Connections 60% $\pm$ 27% 62% $\pm$ 28% 65% $\pm$ 29%
@@
Group 1 (Frontal/Fronto-Central) showed the highest mean accuracy across all window lengths, followed by Group 2 (Central-Parietal) and Group 3 (Parieto-Occipital). Group 4 (Longitudinal Connections) showed the lowest mean accuracy, though still above chance level.
@@ 3.5 Frequency Band Analysis
Figure 5: Violin plots showing the distribution of classification accuracy for each electrode triplet across different frequency bands using 10-second windows.
3.6 Group-Frequency Interactions
@@
1: Frontal/Fronto-Central 56% $\pm$ 25% 59% $\pm$ 27% 62% $\pm$ 28%
2: Central-Parietal 54% $\pm$ 24% 64% $\pm$ 29% 60% $\pm$ 27%
3: Parieto-Occipital 53% $\pm$ 23% 59% $\pm$ 27% 59% $\pm$ 26%
4: Longitudinal Connections 53% $\pm$ 24% 60% $\pm$ 27% 62% $\pm$ 28%
@@
1: Frontal/Fronto-Central 53% $\pm$ 24% 52% $\pm$ 23% 55% $\pm$ 24%
2: Central-Parietal 55% $\pm$ 25% 50% $\pm$ 22% 52% $\pm$ 23%
3: Parieto-Occipital 49% $\pm$ 22% 48% $\pm$ 21% 60% $\pm$ 27%
4: Longitudinal Connections 52% $\pm$ 23% 51% $\pm$ 23% 53% $\pm$ 24%
Table 8: Mean classification accuracy for each electrode group across different frequency bands using 10-second windows. Values represent mean accuracy $\pm$ standard deviation across all electrode triplets within each group.
Each electrode group showed distinct patterns of frequency-specific discriminative power. Group 1 (Frontal/Fronto-Central) showed the highest accuracy in the alpha band. Group 2 (Central-Parietal) showed peak performance in the theta band. Group 3 (Parieto-Occipital) showed notable performance in the gamma band. However, the highest gamma band accuracy was achieved by the central-parietal triplet {CP1, Pz, CP2} from Group 2. Group 4 (Longitudinal Connections) showed highest accuracy in the alpha band.
@@ 3.7 Topographical Patterns
Topographical maps were generated to visualize the spatial distribution of classification accuracy across the scalp for each frequency band. Figure 6 presents the topographical distribution of classification accuracy across different frequency bands using 20-second windows, revealing distinct spatial patterns of discriminative power. These maps demonstrate how the ability to distinguish between PD patients and healthy controls varies across the scalp and frequency spectrum.
@@
It is important to note that the color scales vary across different topographical maps, with yellow representing different peak accuracy values in each map (ranging from approximately 56% in some bands to 74% in others). This calibration allows for better visualization of the spatial patterns within each frequency band. Despite these scale differences, consistent spatial patterns emerge across window lengths, indicating robust regional EEG alterations associated with PD. In particular, the gamma band shows a focal peak of high discriminative power specifically in the central-parietal region, representing the most striking and well-defined spatial pattern across all frequency bands. This midline emphasis in gamma oscillations is particularly notable as it achieves the highest individual frequency-specific accuracy (74% in the {CP1, Pz, CP2} triplet) and persists across all window lengths. The theta band demonstrates bilateral centro-parietal discriminative power. The alpha band shows a more distributed pattern with multiple regions of high discriminative power.
Figure 7: Comparison of gamma band (40-62.4 Hz) topographical maps across different window lengths (5, 10, and 20 seconds). The central-parietal focus of high discriminative power is present across all window lengths.
A consistent pattern across multiple frequency bands is the right-hemisphere emphasis in classification accuracy, particularly evident in the delta, beta2, and gamma bands. This lateralization of EEG alterations suggests asymmetric neurophysiological effects of PD on brain oscillatory activity, with more pronounced or detectable changes in right-hemisphere regions involved in visuospatial processing and sensorimotor integration.
@@ 4.1.1 Functional Group Analysis
Group 1 (Frontal/Fronto-Central regions) achieved the highest overall accuracy (69% with 20-second windows), with the central electrode triplet {C3, CZ, C4} reaching 76% accuracy. This finding aligns with PD’s well-documented disruption of cortico-basal ganglia-thalamo-cortical motor circuits [1, 2]. The central electrodes’ high discriminative power likely reflects underlying pathophysiological changes in primary motor and premotor areas essential for movement planning and execution.
Group 2 (Central-Parietal regions) demonstrated notable frequency-specific patterns, with particularly high classification accuracy in the theta band (64% mean, individual triplets reaching 67%) and, most remarkably, the central-parietal triplet {CP1, Pz, CP2} achieving 74% accuracy in the gamma band. These regions mediate sensorimotor integration—translating sensory input into motor commands—functions increasingly recognized as compromised in PD [19, 20]. The frequency-specific performance suggests distinct oscillatory mechanisms are affected in these integration processes.
Group 3 (Parieto-Occipital regions) showed moderate overall accuracy but interesting frequency-specific performance. The involvement of these visual processing and visuospatial integration regions supports growing evidence that visual dysfunction represents a significant non-motor component of PD pathophysiology [21].
Group 4 (Longitudinal Connections) captured anterior-to-posterior connectivity patterns with notable performance in the alpha band (62%). This suggests PD disrupts long-range communication networks, potentially explaining the integration difficulties observed clinically.

### v2

# Oscillatory Signatures of Parkinson’s Disease: Central and Parietal EEG Alterations Across Multiple Frequency Bands

@@ Abstract
This study investigates EEG as a potential early biomarker by applying deep learning techniques to resting-state EEG recordings from 31 subjects (15 with PD and 16 healthy controls). EEG signals underwent preprocessing to remove tremor artifacts before classification with CNNs using wavelet-based electrode triplet images. Our analysis across different brain regions and frequency bands showed distinct spatial-spectral patterns of PD-related neural oscillations. We identified high classification accuracy (76%) using central electrodes (C3, Cz, C4) with full-spectrum 0.4-62.4 Hz analysis and 74% accuracy in right parietal regions (P8, CP6, P4) with 10-second windows. Bilateral centro-parietal regions showed strong performance (67%) in the theta band (4.0-7.79 Hz), while multiple areas demonstrated some sensitivity (65%) in the alpha band (7.8-15.59 Hz). We also observed a distinctive topographical pattern of gamma band (40-62.4 Hz) alterations specifically localized to central-parietal regions, which remained consistent across different temporal windows. In particular, we observed pronounced right-hemisphere involvement across several frequency bands. Unlike previous studies that achieved higher accuracies by potentially including tremor artifacts, our approach isolates genuine neurophysiological alterations in cortical activity. These findings suggest that specific EEG-based oscillatory patterns, especially in central and parietal regions and across multiple frequency bands, may provide diagnostic information for PD, potentially before the onset of motor symptoms.
@@
Parkinson’s disease , Electroencephalography , Deep learning , Theta oscillations , Alpha oscillations , Central-parietal networks , Biomarkers
@@ 1 Introduction
Parkinson’s disease (PD) is a progressive neurodegenerative disorder characterized primarily by motor symptoms including resting tremor, rigidity, bradykinesia, and postural instability [1]. PD is associated with the loss of dopaminergic neurons in the substantia nigra, resulting in dopamine deficiency [2]. Despite decades of research, current diagnostic approaches still rely heavily on clinical observation of motor symptoms and response to dopaminergic medication—a significant limitation given that these manifestations typically appear only after 60-80% of substantia nigra neurons have already degenerated [3].
@@
The spatial localization of PD-related EEG abnormalities presents another substantial challenge. Yet emerging evidence suggests significant involvement of parietal and occipital cortices in PD pathophysiology, with Tessitore et al. demonstrating altered resting-state functional connectivity in posterior cortical networks that correlates with specific cognitive symptoms [14]. The reliance on predetermined regions of interest in most previous studies has potentially obscured important spatial patterns of EEG abnormalities across the cortex.
@@
The present study addresses these gaps by systematically evaluating EEG activity across multiple brain regions and frequency bands without preconceived notions about which combinations should prove most discriminative. By transforming EEG signals into wavelet-based images and grouping spatially adjacent electrodes, the approach captures local patterns of neural synchronization while addressing the limitations of traditional frequency analysis methods. This methodology enables comprehensive mapping of PD-related EEG alterations across the entire cortex and frequency spectrum.
@@
    To systematically evaluate the discriminative power of different brain regions (electrode groups) and frequency bands without the regional biases evident in previous research, with particular emphasis on central and parietal regions and frequency bands (theta and alpha) that may provide the most discriminative information.
@@
    To identify specific spatial-spectral patterns yielding optimal classification accuracy, with particular attention to central and parietal regions and their frequency-specific alterations that might serve as novel neurophysiological biomarkers for PD.
@@ 2.1 Dataset
It is important to note that the dataset has certain limitations acknowledged by the original data collectors. The Unified Parkinson’s Disease Rating Scale (UPDRS) assessments were conducted by laboratory personnel who had completed online training rather than by board-certified neurologists, which may introduce some uncertainty in clinical evaluations.
Given the heterogeneous presentation of PD across patients and the relatively small sample size, we employed a cross-validation approach to maximize the statistical validity of our findings while acknowledging the limitations inherent to the dataset size.
@@ 2.2 EEG Preprocessing
The EEG data was downsampled from 512 Hz to 128 Hz. This step reduced computational load while maintaining sufficient temporal resolution for frequency analysis up to 64 Hz.
Then, a finite impulse response (FIR) bandpass filter to remove frequency components below 0.4 Hz and above 63.5 Hz. This eliminated slow drifts, DC offset, and high-frequency noise while preserving the frequency bands of interest (delta, theta, alpha, beta, and gamma).
@@ 2.3 Data Preparation
    Increase effective sample size by creating multiple examples from each participant’s recording.
@@
    Compare the classification performance across different temporal scales and assess the stability and consistency of the results.
For each window, we applied continuous wavelet transform (CWT) to convert the one-dimensional time series into two-dimensional time-frequency representations. The resulting wavelet coefficients were analyzed across six frequency bands: delta (0.4-3.99 Hz), theta (4.0-7.79 Hz), alpha (7.8-15.59 Hz), beta1 (15.6-31.19 Hz), beta2 (31.2-39.99 Hz), and gamma (40-62.4 Hz).
To prepare the data for input to the convolutional neural network (CNN), we organized the EEG channels into 20 groups of three spatially adjacent electrodes based on their proximity on the scalp. Figure 1 shows 6 triplets (groups of 3 channels) showing the highest classification accuracy. Altogether these were the 20 electrode triplets split into 4 groups:
Group Electrode Triplets Brain Regions
1 {AF3, Fz, AF4}, {FC1, Fz, FC2}, Frontal and Fronto-Central
{C3, Cz, C4}, {FC5, F3, FC1}, Regions
{FC6, F4, FC2}
2 {CP1, Pz, CP2}, {P3, Pz, P4}, Central-Parietal Regions
{PO3, Pz, PO4}, {CP5, C3, CP1},
{CP6, C4, CP2}
3 {O1, Oz, O2}, {PO3, O1, Oz}, Parieto-Occipital Regions
{PO4, O2, Oz}, {P7, CP5, P3},
{P8, CP6, P4}
4 {F7, T7, P7}, {F3, C3, P3}, Longitudinal/Cross-Regional
{Cz, Pz, Oz}, {F4, C4, P4}, Connections
{F8, T8, P8}
@@ 3.1 Overall Classification Performance
5-second windows 58.4% $\pm$ 29.8%
10-second windows 59.9% $\pm$ 34.4%
20-second windows 60.2% $\pm$ 37.5%
Table 4: Mean classification accuracy across all electrode triplets for different window lengths using the full frequency spectrum. Note that the differences across window lengths are minimal, suggesting that PD-related EEG features are present at various temporal scales.
@@ 3.3 Group-Level Performance
1: Frontal/Fronto-Central 52.8% $\pm$ 31.6% 53.4% $\pm$ 36.0% 52.6% $\pm$ 37.6%
2: Central-Parietal 63.0% $\pm$ 28.6% 64.2% $\pm$ 33.8% 62.8% $\pm$ 37.0%
3: Parieto-Occipital 60.0% $\pm$ 27.8% 62.6% $\pm$ 32.4% 64.2% $\pm$ 36.2%
4: Longitudinal Connections 57.6% $\pm$ 31.0% 59.2% $\pm$ 35.4% 61.2% $\pm$ 39.2%
@@
[…]
@@ 3.4 Frequency Band Analysis
The specificity of theta band sensitivity to bilateral centro-parietal regions is particularly noteworthy, as these regions are known to be involved in sensorimotor integration functions. The central-parietal region ({CP1, Pz, CP2}) from Group 2 showed moderate performance in the gamma band. Overall, the theta and alpha bands demonstrated the most consistent discriminative power across different regions, suggesting their particular relevance for PD classification.
Figure 5: Violin plots showing the distribution of classification accuracy using 10-second windows for each electrode triplet across different frequency bands: delta (top-left), theta (top-right), alpha (mid-left), beta1 (mid-right), beta2 (bottom-left), gamma (bottom-right).
@@
1: Frontal/Fronto-Central 49.0% $\pm$ 18.0% 62.0% $\pm$ 26.0% 64.0% $\pm$ 28.0%
2: Central-Parietal 52.0% $\pm$ 20.7% 63.7% $\pm$ 28.3% 57.7% $\pm$ 27.0%
3: Parieto-Occipital 50.5% $\pm$ 25.0% 61.0% $\pm$ 27.5% 62.5% $\pm$ 25.5%
4: Longitudinal Connections 53.5% $\pm$ 20.5% 63.0% $\pm$ 27.0% 62.0% $\pm$ 26.0%
@@
1: Frontal/Fronto-Central 58.0% $\pm$ 31.0% 49.0% $\pm$ 23.0% 55.0% $\pm$ 36.0%
2: Central-Parietal 60.7% $\pm$ 29.7% 56.7% $\pm$ 25.7% 59.7% $\pm$ 33.7%
3: Parieto-Occipital 54.5% $\pm$ 28.0% 50.0% $\pm$ 24.5% 48.0% $\pm$ 32.0%
4: Longitudinal Connections 57.5% $\pm$ 29.0% 51.0% $\pm$ 23.5% 56.0% $\pm$ 32.5%
Table 8: Mean classification accuracy for each electrode group across different frequency bands using 10-second windows. Values represent mean accuracy $\pm$ standard deviation across electrode triplets within each group.
Each electrode group showed distinct patterns of frequency-specific discriminative power. Group 1 (Frontal/Fronto-Central) showed the highest accuracy in the alpha band (64.0% $\pm$ 28.0%). Group 2 (Central-Parietal) showed peak performance in the theta band (63.7% $\pm$ 28.3%). Group 3 (Parieto-Occipital) showed highest accuracy in the alpha band (62.5% $\pm$ 25.5%). Group 4 (Longitudinal Connections) showed highest accuracy in the theta band (63.0% $\pm$ 27.0%). This regional specificity of frequency band alterations suggests that PD affects different functional networks through distinct oscillatory mechanisms, which may reflect the heterogeneous pathophysiology of the disease.
@@ 3.5 Topographical Patterns
Topographical maps were generated to visualize the spatial distribution of classification accuracy across the scalp for each frequency band. It is important to note that the color scales vary across different topographical maps, with yellow representing different peak accuracy values in each map (ranging from approximately 56% in some bands to 74% in others). This calibration allows for better visualization of the spatial patterns within each frequency band. Despite these scale differences, consistent spatial patterns emerge across window lengths, indicating robust regional EEG alterations associated with PD.
Figure 6 presents the topographical distribution of classification accuracy across different frequency bands using 20-second windows, showing distinct spatial patterns of discriminative power.
@@
A particularly finding from our topographical analysis is the remarkable consistency of gamma band (40-62.4 Hz) spatial distribution across different window lengths (Figure 7). This spatial specificity, combined with its temporal stability across different window lengths (5, 10, and 20 seconds), suggests that gamma band alterations reflect a fundamental aspect of PD pathophysiology rather than a transient feature, that is also supported by [20]. While the central-parietal region ({CP1, Pz, CP2}) showed moderate performance in the gamma band overall, the distinct topographical pattern of gamma alterations compared to other frequency bands indicates that high-frequency neural synchronization in central-parietal regions may represent a specific marker of the disease. This contrasts with lower frequency bands (theta and alpha) which show more distributed patterns of discriminative power across multiple brain regions.
Figure 7: Comparison of gamma band (40-62.4 Hz) topographical maps across different window lengths (5, 10, and 20 seconds).
Further, a consistent pattern across multiple frequency bands is the right-hemisphere emphasis in classification accuracy, particularly evident in the delta, beta2, and alpha bands. This lateralization of EEG alterations suggests asymmetric neurophysiological effects of PD on brain oscillatory activity, with more pronounced or detectable changes in right-hemisphere regions involved in visuospatial processing and sensorimotor integration.
@@ 4.1.1 Functional Group Analysis
Group 1 (Frontal/Fronto-Central regions) showed variability in classification performance, with the central electrode triplet {C3, CZ, C4} significantly outperforming other triplets in this group and reaching 76% accuracy with 20-second windows. This finding aligns with PD’s well-documented disruption of cortico-basal ganglia-thalamo-cortical motor circuits [1, 2]. The central electrodes’ high discriminative power likely reflects underlying pathophysiological changes in primary motor and premotor areas essential for movement planning and execution. The improved performance with longer time windows (from 68% at 10 seconds to 76% at 20 seconds) suggests that longer sampling periods may better capture the consistent abnormalities in motor-related oscillations.
Group 2 (Central-Parietal regions) demonstrated notable frequency-specific patterns, with particularly high classification accuracy in the theta band (64% mean, individual triplets reaching 67%). These regions mediate sensorimotor integration—translating sensory input into motor commands—functions increasingly recognized as compromised in PD [21, 22]. The frequency-specific performance suggests distinct oscillatory mechanisms are affected in these integration processes.
Group 3 (Parieto-Occipital regions) showed strong performance, particularly the right parietal region {P8, CP6, P4} achieving 74% accuracy with 10-second windows. The involvement of parietal regions supports growing evidence of parietal cortex dysfunction in PD [23], potentially contributing to visuospatial processing and sensorimotor integration deficits observed clinically. The comparable performance of this region across 10-second (74%) and 20-second (73%) windows suggests that the discriminative features in this region are relatively stable and can be effectively captured even in shorter time periods.
Group 4 (Longitudinal Connections) captured anterior-to-posterior connectivity patterns with notable performance in both the theta (63%) and alpha (62%) bands. This suggests PD disrupts long-range communication networks, potentially explaining the integration difficulties observed clinically.